OpenStudy (anonymous):

How can I turn 4y = x+3 into slope form ? I have to graph it but I'm confused on how to solve it..

OpenStudy (triciaal):

when x = 0 what is y? when y = 0 what is x?

OpenStudy (triciaal):

you have 4y to get y divide by 4 4y = x + 3 y= (x+3)/4 think of fractions with same denominators y = x/4 +3/4